OpenStudy (anonymous):

PLEASE HELPPPPP The function F(t)=-t^2+2t+3 models the depth of water (in feet)in large grenade ditch, where t is measured in hours and t=0 corresponds to the moment that a summer storm has ended, a. Write F(t) in vertex form by completing the square. b. State and interpret the vertex of y=F(t) as a max. or min. of the real life situation.

OpenStudy (thomas5267):

What is a grenade ditch?

OpenStudy (anonymous):

sorry it's a drainage ditch @thomas5267

OpenStudy (thomas5267):

Do you know how to complete the square?

OpenStudy (anonymous):

not sure can u solve it

OpenStudy (thomas5267):

Suppose the original quadratic equation is \(mx^2+nx+u\), completing the square means that you rewrite the quadratic equation into \(a(x+b)^2+c\). \[ \begin{align*} a(x+b)^2+c&=a(x^2+2bx+b^2)+c\\ &=ax^2+2abx+ab^2+c\\ &=ax^2+2abx+(ab^2+c) \end{align*} \] So \(a=m\), \(2ab=n\) and \(u=ab^2+c\)

OpenStudy (thomas5267):

Do you have any idea what I am talking about?